an eguston For the sfontencons Fissiesn 25 tvo wuck than are idestrual to on a no theR aud Four ea trons
Answer #1

Let the product be X

Balance mass number on both sides

mass of Fm = mass of n + 2*mass of X

1*256 = 4*1 + 2*M

M = 126

Balance atomic number on both sides

atomic of Fm = atomic of n + 2*atomic of X

1*100 = 4*0 + 2*A

A = 50

Atomic number 50 is for element Sn

So, the product is

126-Sn

Answer:

126

Sn

50